OpenStudy (anonymous):

A fence is to be built to enclose a rectangular area of 280 square feet. The fence along three sides is to be made of material that costs 5 dollars per foot, and the material for the fourth side costs 15 dollars per foot. Find the dimensions of the enclosure that is most economical to construct.

OpenStudy (amistre64):

Area = LW Perimeter = L1 + L2 + 2W 180 = LW 2L(5) + W(5) + L(15) = cost

OpenStudy (amistre64):

since 180 = LW, we can say that W = 180/L

OpenStudy (amistre64):

that L(15) up there should be W(15) 2L(5) + (180/L)(5) + (180/L)(15) = Cost

OpenStudy (amistre64):

10L +900/L + 2700/L = cost

OpenStudy (amistre64):

maybe we need to minimise the perimeter :) P = 2L + 2W P = 2L + 2(180/L) P = 2L + 360/L P = 2L^2 + 360 ---------- take the derivative L

OpenStudy (amistre64):

dP = L(2L) - (2L^2+360)(1) ------------------- L(L)

OpenStudy (amistre64):

4L^2 - 2L^2 - 360 2L^2-360 ----------------- = --------- L^2 L^2 I adjusted for my stupidity :)

OpenStudy (amistre64):

when the top is 0 we got a min or a max

OpenStudy (anonymous):

sorry just a little confused because it 280 but it still works i can just go back and plug in my numbers

OpenStudy (amistre64):

2(L^2 -180) = 0 L^2 - 180 = 0 L^2 = 180 L = +sqrt(180)

OpenStudy (amistre64):

...... lol ..........

OpenStudy (amistre64):

i forgot to adjust for all my stupidity lol

OpenStudy (anonymous):

it ok you have it i see just a mix up of numbers is all...

OpenStudy (amistre64):

L = sqrt(280) :)

OpenStudy (amistre64):

W = 280/sqrt(280) W = sqrt(280) as well, so its a square I guess...

OpenStudy (amistre64):

Well, that would maximise the area ..... but it doesnt help with costs :) lets see what costs gets us

OpenStudy (amistre64):

2L(5) + (280/L)(5) + (280/L)(15) = Cost 10L + 900/L + 4200/L = cost

OpenStudy (amistre64):

10L^2 + 900 + 4200 ------------------ = cost L

OpenStudy (amistre64):

stop me when I get lost :)

OpenStudy (amistre64):

10L^2 + 5100 ------------ = cost L L(20L) - 1(10L^2 + 5100) ----------------------- = cost' L^2

OpenStudy (anonymous):

wait i am confused....it wants to know the dimensions...so why are we looking for cost

OpenStudy (amistre64):

20L^2 -10L^2 -5100 = 0 10L^2 - 5100 = 0 10(L^2 -510) = 0 L^2 - 510 = 0 L=sqrt(510) because..... the diminsions dont need to be "minimized for area, they need to be minimized for cost...makes sense?

OpenStudy (anonymous):

yea...

OpenStudy (amistre64):

lets plug sqrt(510) into the calulator to give us a roundabout: 22.58 is what I get for one side :) 280 = LW 280 = 22.58W 280/22.58 = W 12.40 = W

OpenStudy (amistre64):

Now figure out how much we need for the most expensive side :)

OpenStudy (amistre64):

2(22.58)(5) + 12.40(5) + 12.40(15) =? 2(12.40)(5) + 22.58(5) + 22.58(15)? which is cheaper?

OpenStudy (anonymous):

my answer are not correct....thats what my website i have to submit my answer into is telling me...so i don't know if there was a mix up somewhere

OpenStudy (amistre64):

maybe..... I could of thought it wrong as well. but to me, to minimize cost, we would derive the equation for Price and not Area, figure out what a side would be for that price then figure the rest out...

OpenStudy (amistre64):

280/L or 280/W doesnt matter, since its just a name for a side..right?

OpenStudy (amistre64):

polpak is here.... hell know what I did wrong ;)

OpenStudy (amistre64):

hit "post" lol

OpenStudy (anonymous):

Ok, we have \[Area = LW = 280 \rightarrow W = 280/L\] \[Cost = 5(L + 2W) + 15L \] \[\rightarrow Cost = 20L + (5*2)(280)/L \]

OpenStudy (anonymous):

Thanks figured it out...

OpenStudy (anonymous):

\[\frac{d}{dL}Cost = 20 - \frac{2800}{L^2}\] The cost is a min when the derivative is 0 \[\rightarrow 0 = 20-\frac{2800}{L^2}\]

OpenStudy (anonymous):

Ok cool.

OpenStudy (amistre64):

so where did I miss it?

OpenStudy (anonymous):

I can't tell, but you have a lot of cost functions and none of them look the same as the one I have.

OpenStudy (amistre64):

:) they are similar.... I was just going over that to see if I made a mistake

OpenStudy (amistre64):

I did 2L(5) + 180(5)/L + 280(15)/L for cost... but that should be correct either way....

OpenStudy (amistre64):

make that 280 up there

OpenStudy (anonymous):

If those are both 280 instead of 180 it should be ok.

OpenStudy (anonymous):

It just means that your L is my W.

OpenStudy (amistre64):

it is...they are :)

OpenStudy (amistre64):

10L +900/L + 4200/L = cost

OpenStudy (anonymous):

Ok then you would have \[Cost' = 10 -900/L^2 - 4200/L^2\]

OpenStudy (amistre64):

right... you derived them seperately which is fine :)

OpenStudy (amistre64):

10L^2 - 5100 = 0 L^2 - 510 = 0

OpenStudy (amistre64):

L=sqrt(510).... is it right? the same as yours?

OpenStudy (amistre64):

you get L = sqrt(140)

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

But my L is your W remember.

OpenStudy (amistre64):

hmmm....... maybe thats right, havent checked that out yet :)

OpenStudy (amistre64):

11.83 = you 12.40 = me.....

OpenStudy (anonymous):

Hrm. Lemme check that cost again.

OpenStudy (amistre64):

280*5 = 1400 ... not 900

OpenStudy (anonymous):

Ah. That'll do it.

OpenStudy (amistre64):

sqrt(560).... might be better :)

OpenStudy (amistre64):

Yep... thats it.... I just forgot how to multiply integers LOL

OpenStudy (anonymous):

Yep, much better.

OpenStudy (anonymous):

Thats why I always save the plugging the numbers part for last.

OpenStudy (anonymous):

Or as late as I can without making it look horrible.

OpenStudy (amistre64):

yeah.... itda helped if I started out with the 280 and worked it, but..... you know :)